The sum of 3 numbers is 137. The second number is 4 more than 2 times the first number. The third is 5 less than three times the first number. Find the 3 numbers
Answer by venugopalramana(3286)

The sum of 3 numbers is 137. The second number is 4 more than 2 times the first number. The third is 5 less than three times the first number. Find the 3 numbers
let the 1st.number=x
two times this number=2x..and 4 more than this =2x+4=second number.
three times the first number =3x...and 5 less than this =3x-5=third number
sum of the 3 numbers=x+2x+4+3x-5=137
6x=137-4+5=138
x=138/6=23
so the three numbers are 23,2*23+4=50,and 3*23-5=64
23,50 and 64 is the answer.